URGENT QUESTION...
in the following sequence of numbers a, ar, ar^2, ar^3...
a= 1st term (t1)
fina an expression in terms of a,r for
i) 8 th term (t8)
ii) 11 th term (t11)
iii) in the sequence 3,6,12..
find the value of the 15th term.

Answer by Alan3354(69443)

in the following sequence of numbers a, ar, ar^2, ar^3...
a= 1st term (t1)
= at^0
term t = ar^(t-1)
-------------------
find an expression in terms of a,r for
i) 8 th term (t8)
= ar^7
ii) 11 th term (t11)
= ar^10
iii) in the sequence 3,6,12..
find the value of the 15th term.
r = 2
t1 = 3 = 3*r^0
t15 = 3*2^14 = 3*16384
t15 = 49152